# Parameterized Prompt for Incremental Object Detection

### v2

@@ 8.1 Single-step Comparison on MS COCO dataset
We quantitatively analyze the impact of $\lambda$, which controls the sparsity of parameterized prompts structure. A larger $\lambda$ enforces sparser weights, and vice versa. We conduct experiments in the PASCAL VOC2007 under the 5+5+5+5 setting. Tab. 7 presents the accuracy results across different $\lambda$ values. We observe that when $\lambda$ is too small (e.g., $1\times 10^{-6}$), insufficient sparsity fails to concentrate critical knowledge in important parameters, leading to poor performance. Conversely, when $\lambda$ is too large (e.g., $1\times 10^{-4}$), excessive sparsity limits the capacity for preservation of task knowledge, leading to performance degradation. $\lambda=1\times 10^{-5}$ strikes the optimal balance, achieving the highest accuracy of 76.4% and providing 1.1% improvement over the baseline without sparse loss. The experiment demonstrates that appropriate sparsity in parameterized prompts can help preserve important knowledge.
@@
$\theta_{init}$ initialized parameterized prompt before training
@@
Table 8: Lookup table for variable definition in the paper.
9 More Explanations
@@
9.1 Parameterized Prompt Fusion

### v4

@@ B.3 Impact of $\lambda$ in Sparse Loss
We quantitatively analyze the impact of $\lambda$, which controls the sparsity of parameterized prompts structure. A larger $\lambda$ enforces sparser weights. We conduct experiments in the PASCAL VOC2007 under the 5+5+5+5 setting. Tab. 7 presents the accuracy results across different $\lambda$ values. We observe that when $\lambda$ is too small (e.g., $1\times 10^{-6}$), insufficient sparsity fails to concentrate critical knowledge in important parameters, leading to poor performance. Conversely, when $\lambda$ is too large (e.g., $1\times 10^{-4}$), excessive sparsity limits the capacity for preservation of task knowledge, leading to performance degradation. $\lambda=1\times 10^{-5}$ strikes the optimal balance, achieving the highest accuracy of 76.4% and providing 1.1% improvement over the baseline without sparse loss. The experiment demonstrates that appropriate sparsity in parameterized prompt structure can help preserve knowledge.
B.4 Proposal Compression
MD-DETR [2] proposes compressing object-related proposals and introduces a localized query retrieval (QL) method. QL uses a fully connected layer with an input dimension of $N\times D$ and an output dimension of $N$ to generate weights for each proposal. At the same time, QL utilizes the assignment obtained by the Hungarian matching criterion as a supervision signal and constructs a cross-entropy regularization term to supervise the weights of each proposal. However, this design leads to a fully connected layer with a parameter size of $N^{2}\times D$, significantly larger than the prompts pool size. The excessive number of parameters not only reduces storage efficiency but also leads to training difficulties. In the code released by MD-DETR, we observe that although the loss in QL does not converge, QL still has a positive effect on accuracy. Inspired by this observation, we designed a Self-Attention Compression (SAC) module to more effectively compress object-related proposals. We first aggregate the feature information of each proposal by applying global average and max pooling operations along the embedding dimension $D$ of the proposal $P$, generating two different proposal descriptors: $F_{avg}$ and $F_{max}$, which represent the average-pooled and max-pooled proposals, respectively. We then apply the shared fully connected layer to both descriptors, fuse the resulting feature vectors through element-wise summation, and apply a sigmoid activation function to obtain the final attention map $M\in\mathbb{R}^{N}$. In short, the attention map is computed as follows:
$M=\sigma(FC(AvgPool(P))+FC(MaxPool(P)))$ (10)
where $\sigma$ denotes the sigmoid activation function, and $FC$ represents the shared fully connected layer. Compared to the $N^{2}\times D$ parameter size required by QL, the fully connected layer in our method contains only $N^{2}$ parameters, significantly reducing the required number of parameters. To verify the effectiveness of SAC, we first integrate it into MD-DETR [2]. We conduct experiments on the PASCAL VOC2007 dataset using the MS COCO pretrained detector. As shown in Tab. 8, replacing QL with our proposed SAC module improves performance, demonstrating that our method achieves more effective proposal compression with fewer parameters.
Table 8: Results ($A{P_{50}}$, %) on different structrues for proposal compression on PASCAL VOC2007 under the 5+5+5+5 setting.
Method 5+5+5+5
1-5 6-20 1-20
MD-DETR - QL 55.2 63.6 61.5
MD-DETR - SAC 68.9 62.3 64.0
MD-DETR - average 39.0 45.0 43.5
P2IOD - SAC 40.9 72.6 64.7
P2IOD - w/o compression 45.9 70.5 64.3
P2IOD - average 74.0 77.2 76.4
Table 9: Lookup table for variable definition in the paper.
@@ B.4 Proposal Compression
We integrate the proposed SAC module into P2IOD, and an inconsistency arises compared to its integration into MD-DETR. As shown in Tab. 8, adding the SAC module does not improve accuracy as in MD-DETR but instead reduces performance. Averaging all background and object proposals achieves the best performance. We attribute this discrepancy to differences in prompt structure. In MD-DETR, the prompt matching mechanism in the prompts pool requires query features to be task-discriminative, so removing background information enables better matching. Additionally, the prompts pool stores a limited number of prompts and linearly combines them. The limited representational capacity forces the prompts pool to retain only the most distinctive information, namely, object-related information. In contrast, the parameterized prompts in P2IOD act as a mapping from the proposal domain to the prompt domain, offering a larger representational space that allows for the inclusion of background information. In detection tasks, including background information helps the detector distinguish between foreground and background. Therefore, removing background information via the SAC module results in suboptimal prompts. The above experiments suggest that in prompt-based IOD, background-related knowledge should be incorporated into the prompts. The prompts pool is not suitable for IOD due to its structural limitations in providing background knowledge. Furthermore, as shown in Tab. 8, removing the averaging operation leads to performance degradation, indicating that proposal compression is necessary. Without proposal compression, the number of parameters in the parameterized prompts increases dramatically. Although this setup preserves complete proposal information, the excessive number of trainable parameters hinders its convergence. The averaging operation maintains a reasonable parameter size while simultaneously retaining both background and object information. Therefore, P2IOD adopts averaging as its proposal compression strategy.
Appendix C More Explanations
@@
C.1 Parameterized Prompt Fusion